OpenStudy (anonymous):

Traveling at 65 miles/hour how many feet can you travel in 22 minutes? (1miles=5280 feet)

OpenStudy (aripotta):

this will tell you feet per minute \[\frac{ 65mi }{ 1hr }x \frac{ 1hr }{ 60\min }x \frac{ 5280ft }{ 1mi }\] once you get feet per minute, just multiply it by 22 to get your answer

OpenStudy (anonymous):

Thanks

OpenStudy (aripotta):

no problemo :)